# Prompting is Not All You Need! Evaluating LLM Agent Simulation Methodologies with Real-World Online Customer Behavior Data

### v6

@@ 4.5 Ablation Study
To evaluate the impact of synthesized reasoning traces on the model’s action generation capability, we conducted an ablation experiment. In the “w/o reasoning” settings, we removed the reasoning from the model input, and we trained and evaluated the model using only the context and action. This experiment helps us to learn whether learning from synthesized reasoning traces improves the model’s ability to generate user actions.
From Table 1, most models exhibited a significant drop in action generation accuracy when trained without reasoning traces, highlighting the importance of explicit reasoning in guiding predictions. A similar decline was observed in final outcome prediction, with F1 scores consistently decreasing across models. For instance, Qwen2.5-7B achieved a 33.86% F1 score with reasoning traces but dropped to 26.92% without, demonstrating that reasoning traces play a crucial role in improving model performance. These results confirm that incorporating synthesized reasoning traces enhances both action generation and final outcome modeling, reinforcing their value in fine-tuning for human behavior simulation.
@@ 4.6 Error Analysis
We analyze model behavior across different error types, comparing two ICL models, Claude, representing general-purpose chat models, and DeepSeek-R1, representing reasoning-augmented models, with a fine-tuned compact model, Qwen2.5–7B. We focus on five distinct error types. Didn’t terminate, didn’t click, and didn’t search indicate cases where the model chose a different action instead of terminating, clicking, or searching, as the human user did. Searched wrong keyword refers to instances where the model performed a search like the user, but used a different (and incorrect) search query. Clicked wrong button captures cases where the model clicked, but on a different button than the one selected by the human user. Illegal actions generated by models are excluded from this analysis. Overall, Claude and DeepSeek-R1 exhibit very similar error profiles, suggesting that the reinforcement learning process used to incorporate reasoning introduces similar biases to those seen in standard chat-based ICL models.
A major shared failure mode of ICL models is their tendency to continue sessions even when human users would have terminated them by closing the browser window. This aligns with earlier findings that ICL agents are more likely to complete purchases, likely because current LLMs are optimized to fulfill task goals rather than follow subtle social cues or termination heuristics. In contrast, Qwen2.5–7B demonstrates more accurate alignment with human termination behavior.
We also find that Qwen2.5–7B better captures iterative search behavior commonly seen in human users, such as retrying search with a different search keyword after an unsatisfactory result or correcting typos in queries. In comparison, ICL models often stick to the original search result and conduct another type of action when a real human user would have made another search. These findings suggest that fine-tuning not only reduces the overall error rate but also enhances the model’s ability to capture fine-grained user behavior patterns, such as correcting typos, retrying failed searches, and deciding when to terminate a session—behaviors commonly observed in real users but often overlooked by ICL models.
@@ 5.1 Action Misalign Between Human and LLM Agents
Existing research has shown that large language models (LLMs) can generate highly “believable” human behavior simulations, supporting various interactive and social simulation scenarios. However, our results indicate that general-purpose pretrained LLMs, despite their strong ability in a variety of tasks and applications, struggle to accurately generate user actions. This is reflected in their performance—DeepSeek-R1 and Claude 3.5 Sonnet v2, for example, achieved only 11.86% and 11.69% accuracy, respectively, on action generation. These findings suggest that while LLMs may appear human-like in language-based tasks and produce plausible and believable outputs, accurate behavioral prediction requires additional fine-tuning and explicit alignment with real human actions.
Building on this observation, we find that current LLM agents are often behaviorally misaligned with real users, particularly in the shopping domain. For example, as shown in Figure 2 and Figure 3, out-of-the-box LLMs tend to make more purchases than human users. We hypothesize that this stems from the training objectives of many LLM agents, which are typically optimized for task completion (e.g., sending emails) and evaluated based on completion rates and task efficiency. To enable accurate human behavior simulation, it is essential to close this behavioral gap and ensure models capture not just goal-oriented actions, but also the nuanced, often inefficient behaviors of real users.
The ability to simulate nuanced user behavior, including exploratory actions and natural errors such as typos, opens up exciting new research directions. For example, in automatic UI/UX evaluation Lu et al. (2025b), simulating realistic user behavior is crucial: to evaluate a typo-correction feature, an LLM agent must be capable of making typos in a human-like way. Such behavioral fidelity is difficult to achieve without training models on fine-grained, human-aligned data.
@@ 5.3 Future Works
While our study demonstrates the effectiveness of fine-tuning LLMs with reasoning traces for human behavior simulation, several future directions remain. Scaling up the dataset could improve robustness and generalization. Integrating user personas may enable more personalized simulations, and reinforcement learning could be used to refine reasoning generation. Finally, incorporating vision-language models (VLMs) may enhance the model’s understanding of graphical interfaces in complex web environments.
@@ 6 Conclusion
In this work, we present the first quantitative, process-centric evaluation of state-of-the-art large language models (LLMs) for simulating human behavior in an online shopping task. Our study demonstrates that fine-tuning LLMs with real-world human behavioral data and synthesized reasoning traces significantly enhances their ability to generate user actions. Compared to in-context learning and pretrained baselines, our fine-tuned models achieved substantial improvements, including a 13.98% absolute gain in F1 score for final outcome prediction and a 5.4% increase in action generation accuracy.
@@ Appendix A Prompts
Reasoning Synthesize Prompt: {minted}[breaklines]text You will be given a customer’s shopping journey on one of the largest e-commerce platforms globally. you will be given the context (what the user is looking at), the action (what the user did), and your job is to predict the user’s rationale for the action. The rationale should follow Here is an example: example For each action in the input, output a rationale. If the action is "terminate", it means that you didn’t find any desired product and you decided to leave the website by closing the browser window.
@@
[breaklines]markdown <IMPORTANT> Your task is to predict the next action and provide rationale for the action based on the previous actions and context. You need to pretend that you are a user, browsing one of the largest e-commerce platforms globally and searching for a product to purchase. The history action (with details described below) and context will be provided to you. You need to predict the next action and provide rationale for the action. </IMPORTANT>
# Action Space
An action is represented in JSON format, and there are four primary types of actions:
#### 1. ‘type_and_submit‘: Type text into an input field and immediately submit the form. Equivalent to typing text into an input and pressing enter key.
"type": "type_and_submit", "name": "input_name", "text": "search_text"
#### 2. ‘click‘: Click on a button or clickable element identified by ‘name‘.
"type": "click", "name": "clickable_name"
#### 3. ‘terminate‘: When you are unsatisfied with the current search result and you don’t want to buy anything, use ‘terminate‘ to indicate that you want to close the browser window and terminate the task.
"type": "terminate"
# Context Your context will be an **simplified version** of the raw HTML of the one of the largest e-commerce platforms globally page you are looking at. Some interactable elements will be added a unique "name" attribute, which you can use to identify the element to interact with (click or type_and_submit).
# Rationale
The rationale is a first-person sentence of what you are thinking when you make the action. It should be a short sentence that explains why you are making the action.
# Output Format
You need to predict the next action and provide rationale for the action. Your output should follow a strict JSON form:
"action": // action goes here "type": "<type>", … , "rationale": "<rationale>" // rationale goes here, a string
<IMPORTANT> OUTPUT A SINGLE JSON OBJECT, NOTHING ELSE. </IMPORTANT>
Appendix B Example Context
[breaklines,breakafter=_<>]htmlexample_data.html

### v7

@@ 4.5 Ablation Study
To evaluate the impact of synthesized reasoning traces on the model’s action generation capability, we conducted an ablation experiment. In the base setting, we removed reasoning traces from both the training and evaluation stages to isolate the contribution of reasoning-augmented learning. This setup allows us to directly assess whether exposure to synthesized reasoning traces improves the model’s ability to generate user actions and predict outcomes.
From Table 1, most models exhibited substantial improvements in action generation accuracy when trained with reasoning traces, with relative gains ranging from 3.54% to 69.69%. Similar benefits were observed for final outcome prediction, where F1 scores increased by 25.78% to over 600% across models. For example, Qwen2.5-7B achieved a 33.86% F1 score with reasoning traces but dropped to 26.92% without them, underscoring the importance of explicit reasoning guidance. These results confirm that incorporating synthesized reasoning traces enhances both step-wise action generation and final outcome modeling, reinforcing their value in fine-tuning models for human behavior simulation.
@@ 4.6 Error Analysis
We analyze model behavior across different error types, comparing two ICL models, Claude, representing general-purpose chat models, and DeepSeek-R1, representing reasoning-augmented models, with a fine-tuned compact model, Qwen2.5–7B. We focus on five distinct error types. Didn’t terminate, didn’t click, and didn’t search indicate cases where the model chose a different action instead of terminating, clicking, or searching, as the human user did. Searched wrong keyword refers to instances where the model performed a search like the user, but used a different (and incorrect) search query. Clicked wrong button captures cases where the model clicked, but on a different button than the one selected by the human user. Illegal actions generated by models are excluded from this analysis. Overall, Claude and DeepSeek-R1 exhibit very similar error profiles, suggesting that the reinforcement learning process used to incorporate reasoning introduces similar biases to those seen in standard chat-based ICL models. A major shared failure mode of ICL models is their tendency to continue sessions or make the purchase even when human users would have terminated them by closing the browser window. This aligns with earlier findings that ICL agents are more likely to complete purchases, likely because current LLMs are optimized to fulfill task goals rather than follow subtle social cues or termination heuristics. In contrast, fientuned Qwen2.5–7B demonstrates more accurate alignment with human termination behavior.
We also observe that Qwen-2.5–7B better captures iterative search behaviors commonly exhibited by real users, such as retrying a query with corrected keywords or fixing typos after an unsatisfactory result. In contrast, ICL-based models like Claude tend to persist with the original query result and proceed to a different type of action, rather than issuing a refined search. As shown in Table 2, Qwen-2.5–7B more closely matches human revisions in both examples, whereas Claude chooses to visit the product shown on the current search result page. These findings suggest that fine-tuning not only reduces the overall error rate but also enhances the model’s ability to capture fine-grained user behavior patterns, such as correcting typos, retrying failed searches, and deciding when to terminate a session.
@@ 5.1 Action Misalign Between Human and LLM Agents
Existing research has shown that large language models (LLMs) can generate highly “believable” human behavior simulations, supporting various interactive and social simulation scenarios. However, our results indicate that general-purpose pretrained LLMs, despite their strong ability in a variety of tasks and applications, struggle to accurately generate user actions. This is reflected in their performance—DeepSeek-R1 and Claude 3.5 Sonnet v2, for example, achieved only 11.86% and 11.69% accuracy, respectively, on the next action generation task. These findings suggest that while LLMs may appear human-like and produce plausible and believable behaviors, accurate step-wise next action prediction requires additional fine-tuning and explicit alignment with real human actions.
Building on this observation, we find that current LLM agents are often misaligned with real human users, particularly in the online shopping domain. For example, as shown in Figure 2 and Figure 3, out-of-the-box LLMs tend to make more purchases than human users. We hypothesize that this stems from the training objectives of many LLMs, which are typically optimized for the final task completion (e.g., sending emails or making purchases) and evaluated based on task completion and task efficiency (i.e., completing the task with a minimal number of steps). To enable accurate human behavior simulation, it is essential to close this gap and ensure models capture what a real user will do under certain circumstances.
We also advocate for the development of evaluation metrics that reflect the irrational nature of human behavior. On one hand, many user actions are rational and goal-directed, where a single correct next step exists and can be evaluated against a gold standard. On the other hand, users also exhibit irrational or weakly deterministic behaviors, such as semi-random clicks, and typographical errors that even the same person may not replicate consistently. A robust metric should not penalize plausible variants that remain consistent with human trajectories and should explicitly assess an agent’s ability to reproduce human-like error patterns. For example, the capacity to generate human-like typos is also essential for applications like automatic UI/UX evaluation Lu et al. (2025b), where assessing typo correction features requires agents to reproduce the kinds of errors real users make. A human-aligned evaluation should therefore distinguish true errors from acceptable variability and report performance across both rational and non-rational action regimes.
@@ 5.3 Future Works
Future research should move beyond treating reasoning traces as static supervision and instead focus on actively enhancing the model’s reasoning abilities. One direction is to use reinforcement learning to improve the reasoning for action prediction accuracy, where better reasoning is defined by its ability to produce more accurate action prediction. Another direction is to train models to generate reasoning that more closely reflects human cognitive processes, so the reasoning itself can serve as a source of qualitative insight when analyzing or interacting with LLM agents. In addition to advancing reasoning, future work could scale up the dataset to improve robustness, incorporate user personas for more personalized simulations, and incorporate user intents with intent detection methods. Finally, incorporating vision-language models (VLMs) may enhance the model’s understanding of graphical interfaces in complex web environments.
@@ 6 Conclusion
In this work, we present the first quantitative, process-centric evaluation of state-of-the-art large language models (LLMs) for simulating human behavior in an online shopping task. Our study demonstrates that LLMs fine-tuned with real-world human behavioral data and synthesized reasoning traces showed a significant enhancement in their ability to generate user actions. Compared to in-context learning and pretrained baselines, our method achieved substantial improvements, including a 13.98% absolute gain in F1 score for final outcome prediction and a 5.4% increase in action generation accuracy.
@@ Appendix A Prompts
Reasoning Synthesize Prompt: